# DistALANER: Distantly Supervised Active Learning Augmented Named Entity Recognition in the Open Source Software Ecosystem

### v3

@@ Abstract
With the AI revolution in place, the trend for building automated systems to support professionals in different domains such as the open source software systems, healthcare systems, banking systems, transportation systems and many others have become increasingly prominent. A crucial requirement in the automation of support tools for such systems is the early identification of named entities, which serves as a foundation for developing specialized functionalities. However, due to the specific nature of each domain, different technical terminologies and specialized languages, expert annotation of available data becomes expensive and challenging. In light of these challenges, this paper proposes a novel named entity recognition (NER) technique specifically tailored for the open-source software systems. Our approach aims to address the scarcity of annotated software data by employing a comprehensive two-step distantly supervised annotation process. This process strategically leverages language heuristics, unique lookup tables, external knowledge sources, and an active learning approach. By harnessing these powerful techniques, we not only enhance model performance but also effectively mitigate the limitations associated with cost and the scarcity of expert annotators. It is noteworthy that our model significantly outperforms the state-of-the-art LLMs by a substantial margin. We also show the effectiveness of NER in the downstream task of relation extraction  [^1] https://anonymous.4open.science/r/DistALANER-DC9D/README.md.
@@ 1 Introduction
Traditional named entity recognition (NER) models exhibit certain limitations, particularly when dealing with domain-specific data. Primarily, this stems from the fact that NER models are conventionally trained on generic corpora, rendering them less effective when encountering text sourced from specialized domains such as software, legal Trias et al. (2021), biomedical Jensen et al. (2006); Krallinger et al. (2015); Nédellec et al. (2013) or engineering fields Li et al. (2022), which inherently possess distinctive vocabularies and entities. For instance, the word “windows” could denote a well-known operating system in the realm of software and technology, yet simultaneously refer to a commonplace architectural feature in the context of residential or commercial structures. Adapting a generic NER model to handle tasks specific to a certain field often requires extra, specialized training data from that area. However, gathering and annotating this additional data can be a costly and time-consuming task.
@@
Distant supervision Liang et al. (2020); Fang et al. (2021) methods help solve the problem of insufficient labels by automatically generating labeled data for entity recognition. Using a raw text and a dictionary, these methods label entities through exact string matching, then use this data to train advanced neural models for recognizing entities. However, two key challenges arise from this approach. The first is incomplete annotations Jie et al. (2019). Many dictionaries do not fully cover domain-specific entities, leading to a lot of unmatched entities and false-negative labels. Earlier attempts to increase labeled entities involved expanding the dictionary with set rules Liu et al. (2019), but these rules are often hard to apply in other fields. The second challenge is the struggle to identify new, unannotated entities. Even models that are manually trained have difficulty with this due to their limited capabilities.
@@
In the area of open source software development, the need for NER has become increasingly critical Tang et al. (2023). NER plays a pivotal role in deciphering and categorizing textual information into predefined entities such as individual contributors, programming languages, software tools, and project specifications found in software documentation, source code, bug reports, or community discussions Wang et al. (2022). The understanding and categorization of such entities offer deep insights, allowing for effective communication, resource allocation, and decision-making within the open source community. In addition, it assists in knowledge extraction Perera et al. (2020), facilitating a greater comprehension of the system’s functionality Shrimal et al. (2022), feature requirements, software dependencies, or potential bugs Zhou et al. , thereby augmenting the efficiency of software development and maintenance. Further, NER can aid in community management tasks, such as identifying contributors and their areas of expertise or mapping the interactions within the developer community. Consequently, the integration of NER in the open source software domain could dramatically streamline processes, enhance collaboration, and eventually improve the overall quality of the software produced.
@@
Recently, Large Language Models (LLMs) can be highly effective in the software domain Ozkaya (2023). With their ability to understand complex patterns and generate human-like text, they can assist in identifying and classifying key entities such as specific coding languages, software tools, packages, peripherals, or developers mentioned in various sources like code, software documentation, and community discussions. However, LLMs can sometimes prove to be a bottleneck in identifying NER due to security issues, cost and lack of contextual knowledge Banerjee et al. (2024). These limitations arise due to their broad contextual learning from vast corpora that often spans numerous domains, making them less specialized for a particular field like software development. These models might struggle to identify and classify domain-specific entities accurately, given their generic training Qin et al. (2023). Moreover, as LLMs learn from data available up to their last training checkpoint, they may not be aware of new terms or entities introduced in the domain post-training. Taking into account the aforementioned limitations Yang et al. (2023), we put forth an innovative framework, explicitly designed for the software domain trained through bug data, manuals and CQAs. Our objective is to incorporate lightweight models into our framework, allowing them to seamlessly integrate with any system. Our approach involves limited human intervention. By including a variety of methods depicted in Figure 2, we enhance the efficiency of deep learning models. The key contributions of this paper are as follows.
@@
  • In terms of datasets we release the following items - (a) a huge open source domain corpus with both human and system annotations for nine different named entity types (see examples in Figure 1), (b) For each entity type, a large unique lookup table contains relevant entities automatically collected from various sources since 2004 (to check detailed sources, see section 4) and (c) a huge corpus of human annotated entity relation pairs in software domain for the downstream application task [^2] https://zenodo.org/record/8075578.
@@
  • We conduct extensive experiments on two massive datasets with DistALANER achieving the best performance using dictionaries only and no human efforts. On several datasets, our approach is even better than the supervised models and LLMs like GPT-3.5-Turbo [^3] https://openai.com/blog/chatgpt, GPT-$4$https://platform.openai.com/playground and Google-BARD [^5] https://bard.google.com/.
  • In order to demonstrate the effectiveness of the named entities extracted by our method we use it to perform the closely associated problem of relation extraction reporting superior performance.
@@ 2 Related work
Named entity recognition in software ecosystems: Initially Ye et al. (2016) proposed S-NER for identifying software-related entities, focusing on high-level entity types such as programming language and API, amongst others. They used a subset of the 2015 StackOverflow dataset and manually identified software-specific entities. Human annotators labeled additional posts, with the annotated data serving for supervised learning and model testing. Li et al. (2020); Zhou et al. introduced a unified framework employing machine reading comprehension for nested and flat Named Entity Recognition (NER). The authors devised a technique to extract the answer span for an entity, instead of labeling sequences. Extensive experimentation was carried out on nested and flat NER datasets to verify their proposed model. Tabassum et al. (2020) presented SoftNER for recognizing code tokens and software-related entities in natural language and programming texts. They used a programming-related dataset from StackOverflow, introducing an NER corpus for the programming domain. This corpus, annotated with 20 self-defined entity types, was used to train BERT Devlin et al. (2019), showing improvement over a pretrained model. In another study Kocaman and Talby (2020), the authors presented BiLSTM-CNN-Char, a deep learning architecture for extracting entities from biomedical notes and reports. The architecture was claimed to be efficient without relying on heavy transformer-based contextual embeddings. Recently, few-shot, multigrained, and nested NER have gained attention Ding et al. (2021). A notable contribution is the FEW-NERD dataset, a large-scale human-annotated collection featuring two-level entity type annotations. The authors in Xia et al. (2019) proposed Multi-Grained NER (MGNER), capable of identifying named entities at multiple granularities. The framework can handle non-overlapping or nested entity mentions and consists of a detector and a classifier component for efficient entity recognition.
Distant supervision based named entity recognition: In Wang et al. (2019); Fang et al. (2021) the authors introduced a novel method of dictionary extension for extracting new entities utilizing a type expanded model. This method’s primary aim was to enhance the performance of distantly supervised NER methods in automating data labeling and entity identification. The evaluation of their approach on diverse datasets demonstrated superior performance in comparison to the existing state-of-the-art distantly supervised systems. Another work by Yu Meng ; Peng et al. (2019) presented a novel method for training NER models using distantly labeled data, typically acquired by associating entity mentions in raw text with corresponding entity types. Liang et al. (2020) put forward a unique computational framework, BOND, which leverages pretrained language models to enhance NER model performance. GPT-NER by Wang et al. (2023) is a first of its kind approach that recasts the sequence labeling task into a generative framework.
Our work introduces a distantly supervised method to auto-annotate named entities in the open-source software systems. This approach is very effective is a low resource setting where gold labels are very limited. We further show that this NER pipeline improves the task of relation extraction.
@@ 3 Dataset
In this paper, we utilize two datasets specifically from the Ubuntu ecosystem: (i) the Ubuntu bug repository and (ii) the Launchpad question-answering repository. We employ the first dataset for training the model, while we use the latter for evaluating its performance. The motivation lies in selecting bug data because it is an ideal resource for training an open-source NER model for several reasons. First, it provides a rich source of diverse and complex natural language text, which includes technical terminology, software components, and descriptions of problems and solutions, making it a well-suited resource for understanding and learning the language structure and context within the open-source domain. Second, bug reports often involve specific named entities such as software component names, version numbers, and user handles, among others, providing a plethora of examples for NER tasks. Further, the nature of bug tracking in open-source projects often involves collaboration and communication between various contributors, yielding a wide variety of linguistic styles and expressions. This diversity enhances the model’s robustness and adaptability. Last, as bug data is openly accessible, it aligns with the open-source ethos of shared knowledge, making it an appropriate dataset for open-source NER model training.
Ubuntu bug repository: In our experiment, we use the repository of bugs collected by Hazra et al. (2021). These bugs are mainly reported on packages, conflicts between Ubuntu and Windows and other related events. The dataset contains approximately 270K bugs along with the metadata such as title of the bug, description of the bug, user who posted the bug, comments and their commenters, creation date of the bug, tags of the bugs. In our work, we mainly use the description of the bugs to obtain the named entities. We filter this raw dataset by excluding (i) those bugs that solely reference another bug, for example, “Automatically imported from Debian bug report #257568" [^6] http://bugs.debian.org/257568 and (ii) those bugs that have very small description size ($<60$ words) or exceedingly large description size ($>400$ words). The bugs with very small description size prohibits obtaining meaningful representations while those with very large size routinely include code logs in massive proportions rather than useful text. The dataset [^7] https://launchpad.net/ statistics are presented in Table 1.
Ubuntu question-answering: We choose Ubuntu question-answering dataset for the purpose of evaluation. We collect queries from Launchpad question-answering, which contains questions on the Ubuntu system and the many problems related to it. Each such question has a title, a body, an asker identity, a posting date and time, a set of answers, the answerer identity and the answer posting time. For our purpose, we annotate a total of 500 question-answer pairs. These question-answer pairs are chosen randomly to ensure an unbiased sampling.
@@ 4 Source details
  • Commands: Commands in structured form come from github [^11] https://github.com/nengz/ShellFusion/tree/main. We collect additional commands using TagMe and add them with our active learning based approach from Wikipedia.
  • Packages: We use all the packages assembled by the authors in Hazra et al. (2021).
@@ 5 Preliminaries
Note that the corpora for our experiments are primarily composed of text drawn from bug repositories and we shall therefore define the notations in terms of this particular corpora. In a bug repository $\mathcal{B}$, each bug is denoted by $b_{i}$. A bug $b_{i}$ is a sequence of words $\{w_{1},w_{2},\cdots,w_{n}\}$. We denote the entity types as the $\emph{E}_{ename}$ where ename is the name of the entity type. We define nine entity types as follows – packages (PKG), operating system (OS), organization (ORG), commands (CMD), errors (ERR), file extension (EXT), peripherals (PRP), software components (SOC), and architecture (ARC). In our setting, an entity may contain single word or multiple words (phrases). The entity span can be defined as $\{w_{i},w_{i+1},\cdots,w_{j-1},w_{j}\}$, where $i$ indicates the starting index, $j$ indicates the ending index, and $i\leq j$. We use the conventional IO tagging (inside-outside) method. Given a sequence of words $\{w_{1},\cdots,w_{i},w_{i+1},\cdots,w_{j},\cdots w_{n}\}$ we mark as {${O,\cdots,I_{ename},I_{ename},\cdots,I_{ename},\cdots,O}$}. An example bug text and its IO tags are shown below. Text: After upgrading to Ubuntu 18.04 and thus from Linux 4.13 to Linux 4.15) the Monitor connected via VGA (through DVI-I) shows a ’No Signal’ message after amdgpu takes over from efifb and turns off. Entity labels: {$O,O,O,I_{OS},I_{OS},O,O,$ $O,I_{OS},I_{OS},O,I_{OS},I_{OS},O,I_{PRP},O,O,I_{PRP},$ $O\cdots$}.
@@ 6.1 Stage 1: Construction and matching of dictionary
At the first stage, we build the dictionary of entities and their respective entity types. To build the dictionary, we use existing knowledge from websites, repositories, and documents of the Ubuntu eco system. For the OS entity, we include all the Ubuntu distributions (obtained from Hazra et al. (2021)) and collect Linux distributions and Windows versions from Wikipedia. For the ARC entity type, we include different writing styles (and versions) of 32 bit and 64 bit (x32 | x64, x386 | amd64, etc.). For CMD type, we collect all the Linux commands from Wikipedia, Ubuntu man pages, and the paper Zhang et al. (2022). Authors in Hazra et al. (2021) collected the list of packages of each Ubuntu distributions (total 20 distributions). For the PKG type, we utilize this list and extract all the unique package names. We build the ERR type by collecting the error codes that usually occur in the Ubuntu system from the Ubuntu wiki page [^14] https://wiki.ubuntu.com/error/and/warning/messages. For the type EXT, we collect data from various Wikipedia pages. For the ORG type, we focus on software organizations and collect the information from Wikipedia. For type PRP, we gather entities from various sources (see section 4 followed by manual inspection. For SOC, we collect data from various unix-related websites as well as Wikipedia pages (See Table 2). Next, we conduct dictionary matching. We consider only exact string matching of the bug description text with the entities corresponding to each entity type. For instance, ‘CurrentDesktop’ is not marked because it contains ‘Desktop’ as a subword. We discard the wrongly identified entities using a set of regex (see Appendix B). Note that, in earlier papers Shang et al. (2018); Wang et al. (2020), researchers have used the AutoPhrase tool  [^15] https://github.com/shangjingbo1226/AutoPhrase to extract phrases from the text, which they later consider for entity identification. We also attempted to apply AutoPhrase to identify entity-containing phrases. However, the tool failed to extract key phrases from our bug descriptions and instead assigns higher scores to irrelevant terms like ‘box’, ‘release’, ‘start’, ‘architecture’, ‘show’, ‘network’, ‘subdevice’, and ‘software’. Hence, we resorted to the above dictionary matching technique.

### v4

@@ Abstract
As the AI revolution unfolds, the push toward automating support systems in diverse professional fields ranging from open-source software to healthcare, and banking to transportation has become more pronounced. Central to the automation of these systems is the early detection of named entities, a task that is foundational yet fraught with challenges due to the need for domain-specific expert annotations amid a backdrop of specialized terminologies, making the process both costly and complex. In response to this challenge, our paper presents an innovative named entity recognition (NER) framework tailored for the open-source software domain. Our method stands out by employing a distantly supervised, two-step annotation process that cleverly exploits language heuristics, bespoke lookup tables, external knowledge bases, and an active learning model. This multifaceted strategy not only elevates model performance but also addresses the critical hurdles of high costs and the dearth of expert annotators. A notable achievement of our approach is its capability to enable pre-large language models (pre-LLMs) to significantly outperform specially designed generic/domain specific LLMs for NER tasks. We also show the effectiveness of NER in the downstream task of relation extraction.
Keywords:
Distant Supervision Active Learning Open Source NER LLM
@@ 1 Introduction
Traditional named entity recognition (NER) models exhibit certain limitations, particularly when dealing with domain-specific data. Primarily, this stems from the fact that NER models are conventionally trained on generic corpora, rendering them less effective when encountering text sourced from specialized domains such as software, legal [25], biomedical [8] or engineering fields [13], which inherently possess distinctive vocabularies and entities. For instance, the word “windows” could denote a well-known operating system in the realm of software and technology, yet simultaneously refer to a commonplace architectural feature in the context of residential or commercial structures. Adapting a generic NER model to handle tasks specific to a certain field often requires extra, specialized training data from that area. However, gathering and annotating this additional data can be a costly and time-consuming task. Distant supervision [15] methods help solve the problem of insufficient labels by automatically generating labeled data for entity recognition. Using a raw text and a dictionary, these methods label entities through exact string matching, then use this data to train advanced neural models for recognizing entities. However, two key challenges arise from this approach. The first is incomplete annotations [9]. Many dictionaries do not fully cover domain-specific entities, leading to a lot of unmatched entities and false-negative labels. Earlier attempts to increase labeled entities involved expanding the dictionary with set rules [16], but these rules are often hard to apply in other fields. The second challenge is the struggle to identify new, unannotated entities. Even models that are manually trained have difficulty with this due to their limited capabilities.
@@
In the area of open source software development, the need for NER has become increasingly critical [24]. NER plays a pivotal role in deciphering and categorizing textual information into predefined entities such as individual contributors, programming languages, software tools, and project specifications found in software documentation, source code, bug reports, or community discussions [28]. The understanding and categorization of such entities offer deep insights, allowing for effective communication, resource allocation, and decision-making within the open source community. Further, NER can aid in community management tasks, such as identifying contributors and their areas of expertise or mapping the interactions within the developer community. Consequently, the integration of NER in the open source software domain could dramatically streamline processes, enhance collaboration, and eventually improve the overall quality of the software produced.
@@
Recently introduced LLMs can be highly effective in the software domain [18]. With their ability to understand complex patterns and generate human-like text, they can assist in identifying and classifying key entities such as specific coding languages, software tools, packages, peripherals, or developers mentioned in various sources like code, software documentation, and community discussions. However, LLMs can sometimes prove to be a bottleneck in identifying NER due to security issues, cost and lack of contextual knowledge. These limitations arise due to their broad contextual learning from vast corpora that often spans numerous domains, making them less specialized for a particular field like software development. These models might struggle to identify and classify domain-specific entities accurately, given their generic training [21]. Moreover, as LLMs learn from data available up to their last training checkpoint; they may not be aware of new terms or entities introduced in the domain post-training. Taking into account the aforementioned limitations [31], we put forth an innovative framework, explicitly designed for the software domain trained through bug data, manuals and CQAs. Our objective is to incorporate lightweight models into our framework, allowing them to seamlessly integrate with any system. Our approach involves limited human intervention. By including a variety of methods depicted in Figure 2, we enhance the efficiency of deep learning models. The key contributions of this paper are as follows.
@@
  • In terms of datasets we release the following items - (a) a huge open source domain corpus with both human and system annotations for nine different named entity types (see examples in Figure 1), (b) For each entity type, a large unique lookup table contains relevant entities automatically collected from various sources since 2004 (to check detailed sources, see section 4) and (c) a huge corpus of human annotated entity relation pairs in software domain for the downstream application task.
@@
  • We conduct extensive experiments on four massive datasets with DistALANER framework achieving the best performance with minimal human efforts. Utilising our framework in conjunction with pre-LLM era models outperforms LLMs like GPT-3.5-Turbo [^1] https://openai.com/blog/chatgpt, GPT-4222https://platform.openai.com/playground, Google-BARD [^3] https://bard.google.com/ and task specialized UniversalNER [^4] https://universal-ner.github.io/ by a substantial margin.
@@ 2 Related work
Named entity recognition in software ecosystems: The landscape of Named Entity Recognition (NER) in software ecosystems has evolved significantly, incorporating diverse approaches and datasets. [32] pioneered S-NER to identify software-related entities like programming languages and APIs using a subset of the 2015 StackOverflow dataset. This effort was complemented by manual annotations for supervised learning. A shift towards a unified framework was made by [14, 34], employing machine reading comprehension to identify entities via answer spans rather than sequence labeling, tested on both nested and flat NER datasets. In parallel, [23] introduced SoftNER, leveraging a programming-related StackOverflow dataset and BERT [2] for enhanced recognition of code tokens and software entities. Meanwhile, [11] showcased BiLSTM-CNN-Char for biomedical entity extraction, emphasizing efficiency without heavy transformer reliance. The field continues to advance with the introduction of the FEW-NERD dataset for few-shot, multigrained, and nested NER [3], alongside [30]’s MGNER framework for multi-granularity entity detection, highlighting the ongoing innovation in entity recognition technologies.
Distant supervision based named entity recognition: In [4, 19], authors introduced innovative methods to enhance distantly supervised Named Entity Recognition (NER) through dictionary extension, type expansion, and leveraging distantly labeled data for model training. These approaches aimed to automate data labeling and entity identification, demonstrating superior performance over existing systems. Additionally, [15] and [27] proposed the BOND framework and GPT-NER respectively, utilizing pretrained language models and a generative framework to further improve NER model efficacy.
@@ 3 Dataset
In this paper, we utilize two datasets specifically from the Ubuntu ecosystem: (i) the Ubuntu bug repository and (ii) software community question-answering repositories. The latter is further subdivided into QAs from three different community posts – Linux, Fedora and Ubuntu thus resulting in a total of four datasets. We use the bug dataset for training the model and the three QA datasets for evaluating its performance. The motivation for selecting bug data for training an open-source NER model are as follows. First, it provides a rich source of diverse and complex natural language text, which includes technical terminology, software components, and descriptions of problems and solutions, making it a well-suited resource for understanding and learning the language structure and context within the open-source domain. Second, bug reports often involve specific named entities such as software component names, version numbers, and user handles, among others, providing a plethora of examples for NER tasks. Further, the nature of bug tracking in open-source projects often involves collaboration and communication between various contributors, yielding a wide variety of linguistic styles and expressions. This diversity enhances the model’s robustness and adaptability. Last, as bug data is openly accessible, it aligns with the open-source ethos of shared knowledge, making it an appropriate dataset for open-source NER model training. Ubuntu bug repository: In our experiment, we use the repository of bugs collected by [5]. These bugs are mainly reported on packages, conflicts between Ubuntu and Windows and other related events. The dataset contains approximately 270K bugs along with the metadata such as title of the bug, description of the bug, user who posted the bug, comments and their commenters, creation date of the bug, tags of the bugs. In our work, we mainly use the description of the bugs to obtain the named entities. We filter this raw dataset by excluding (i) those bugs that solely reference another bug, for example, “Automatically imported from Debian bug report #257568" [^5] http://bugs.debian.org/257568 and (ii) those bugs that have very small description size ($<60$ words) or exceedingly large description size ($>400$ words). The bugs with very small description size prohibits obtaining meaningful representations while those with very large size routinely include code logs in massive proportions rather than useful text. The dataset statistics are presented in Table 1. QA datasets: We choose Ubuntu [^6] https://launchpad.net/ubuntu, Fedora  [^7] https://forums.fedoraforum.org/ and Linux [^8] https://www.linux.org/forums/ question-answering community posts for the purpose of evaluation. These posts contain questions on the respective open-source system and the many problems related to it. Each such question has a title, a body, an asker identity, a posting date and time, a set of answers, the answerer identity and the answer posting time. For our purpose, we annotate a total of 500 question-answer pairs from each community. These question-answer pairs are chosen randomly to ensure an unbiased sampling.
@@ 4 Source details
  • Commands: Commands in structured form come from github [^12] https://github.com/nengz/ShellFusion. We collect additional commands using TagMe and add them with our active learning based approach from Wikipedia.
  • Packages: We use all the packages assembled by the authors in [5].
@@ 5 Preliminaries
[…]
Monitor connected via VGA (through DVI-I) shows a ‘No Signal’ message after
amdgpu takes over from efifb and turns off.
Entity labels: {$O,O,O,I_{OS},I_{OS},O,O,O,I_{OS},I_{OS},O,I_{OS},I_{OS},O,I_{PRP},O,O,I_{PRP},O\cdots$}.
@@ 6.1 Stage 1: Construction and matching of dictionary
At the first stage, we build the dictionary of entities and their respective entity types. To build the dictionary, we use existing knowledge from websites, repositories, and documents of the Ubuntu eco system. For the OS entity, we include all the Ubuntu distributions (obtained from [5]) and collect Linux distributions and Windows versions from Wikipedia. For the ARC entity type, we include different writing styles (and versions) of 32 bit and 64 bit (x32 | x64, x386 | amd64, etc.). For CMD type, we collect all the Linux commands from Wikipedia, Ubuntu man pages. Authors in [5] collected the list of packages of each Ubuntu distributions (total 20 distributions). For the PKG type, we utilize this list and extract all the unique package names. We build the ERR type by collecting the error codes that usually occur in the Ubuntu system from the Ubuntu wikipage [^15] https://wiki.ubuntu.com/error/and/warning/messages.
@@
For the type EXT, we collect data from various Wikipedia pages. For the ORG type, we focus on software organizations and collect the information from Wikipedia. For type PRP, we gather entities from various sources (see section 4) followed by manual inspection. For SOC, we collect data from various unix-related websites as well as Wikipedia pages (See Table 2). Next, we conduct dictionary matching. We consider only exact string matching of the bug description text with the entities corresponding to each entity type. For instance, CurrentDesktop is not marked because it contains Desktop as a subword. We discard the wrongly identified entities using a set of regex (see Appendix). Note that, in earlier paper [29], researchers have used the AutoPhrase tool  [^16] https://github.com/shangjingbo1226/ AutoPhrase to extract phrases from the text, which they later consider for entity identification. We also attempted to apply AutoPhrase to identify entity-containing phrases. However, the tool failed to extract key phrases from our bug descriptions and instead assigns higher scores to irrelevant terms like ‘box’, ‘release’, ‘start’, ‘architecture’, ‘show’, ‘network’, ‘subdevice’, and ‘software’. Hence, we resorted to the above dictionary matching technique.
6.2 Stage 2: Entity distillation and dictionary expansion
In Stage 2, we distill the exactly matched entities and then expand the list of entities for a given entity type using an active learning approach. Active learning enables our system to learn iteratively, refining its understanding as it receives feedback from its interactions. We observe that many phrases get marked as an entity, but in reality, some of them are not entities. Thus, we employ two heuristics to distill the entities: (1) identifying parts-of-speech patterns, and (2) human intervention. Parts-of-speech patterns: Here, we find that the likelihood of a phrase being an entity often depends on the parts-of-speech tag of the words in the phrase or before and after the phrase. For instance, the verb ‘find’ in a sentence is never an entity, even though our dictionary matching may mistakenly identify it as a CMD entity (see more samples in section 7). We adjust the contents of the dictionary through such iterative revisions.